What is the slope m and y-intercept point of the line y = 2x + 1
Found 2 solutions by checkley77, sweetlipstogo:
Answer by checkley77(12844)

Y=Mx+b m=SLOPE & b=THE Y INTERCEPT.
y = 2x + 1
m=2 for the slope & 1 for the Y intercept.

Answer by sweetlipstogo(3)

To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept
y = mx + b
m = 2, b = 1